OpenStudy (welshfella):

Calculate the remainder when 3^293 is divided by 97. I'm not sure how to solve this. My work so far:- 293 = (3*97) + 2 3^293 = 3^(3 * 97 + 2) = 3^2 * (3^3)^97 = 9* 27^97. I'm stuck here.

OpenStudy (anonymous):

is it like a show-your-work problem?

OpenStudy (welshfella):

I guess we need to bring modular arithmetic into this? I find that stuff rather difficult...

OpenStudy (welshfella):

But we are talking remainders so mod arithmetic comes to mind.

OpenStudy (welshfella):

OK i 've found the solution. You use Fermat's Little Theorem

OpenStudy (welshfella):

x^p = x mod p where p is a prime number so here we have 3^293 mod 97 = ( 9*27^97) mod 97 = (9 * 27) mod 97 = (243) mod 97 = 49 mod 97 so the remainder is 49.

OpenStudy (welshfella):

Please check my work

OpenStudy (welshfella):

@ganeshie8

OpenStudy (sparrow2):

i think it's correct

OpenStudy (welshfella):

yea I came across Fermats Little theorem browsing a chapter on modular arithmetic. That was new to me.

OpenStudy (sparrow2):

oh in number theory and arithmetic therma is really important :")and this theorem too

OpenStudy (sparrow2):

there is also thermat's big theorem

OpenStudy (welshfella):

I'd only heard of Fermats Last Theorem whose proof had to wait for centuries!!

OpenStudy (sparrow2):

i think it is thermats big theory you are talking about..it was proveid in 90's wiles needed 7 years to prove :)

OpenStudy (sparrow2):

he wote that he had the proof but didn't have free space to write it on the margin..wiles needed 150 pages or more ;)

OpenStudy (welshfella):

Wow! Thats the one about a^n + b^n= c^n being not true for n = integer >2 right?

OpenStudy (sparrow2):

yeah that is: but it is proved and numbers are found to satisfy that

OpenStudy (welshfella):

There are numbers which satisfy that? That I didn't know. Whole numbers?

ganeshie8 (ganeshie8):

Hey!

OpenStudy (sparrow2):

yaeh,you can google

ganeshie8 (ganeshie8):

Your work looks perfect!

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=3%5E293+mod+97

OpenStudy (welshfella):

Good! I'm beginning to get a better grasp of modular arithmetic. THanks ganeshie.

ganeshie8 (ganeshie8):

Just for the sake of an alternative, notice that : \[x^p\equiv x\pmod{p}\] is same as \[x^{p-1}\equiv 1\pmod{p}\] whenever \(\gcd(x,p)=1\)

ganeshie8 (ganeshie8):

Therefore, rising both sides to \(k\)th power we get : \[(x^{p-1})^k \equiv 1 \pmod{p}\]

OpenStudy (welshfella):

thats looks a useful result

ganeshie8 (ganeshie8):

Since \(97\) is prime, we have : \[(3^{96})^k\equiv 1\pmod{97}\]

ganeshie8 (ganeshie8):

plugging \(k=3\) gives \[3^{288}\equiv 1\pmod{97}\]

ganeshie8 (ganeshie8):

multiply both sides by \(3^5\) and we're done!

OpenStudy (welshfella):

yes 3^5 = 243

OpenStudy (welshfella):

great